OpenStudy (anonymous):

Find the slope of the line whose equation is 4y - 3x + 6 = 0. -3/4 -3/2 3/4

OpenStudy (arabpride):

set the equation to look like: y=mx+b 4y - 3x + 6 = 0 +3x +3x 4y + 6 = 3x -6 -6 4y = 3x - 6 /4 /4 y = 3/4x - 6/4 y = 3/4x - 3/2 since 'm' iz ur slope, then, 3/4 is the value of ur slope.. @hateschoolwithpassion